# Advancing Multi-Organ Disease Care: A Hierarchical Multi-Agent Reinforcement Learning Framework

### v1

@@
Abstract
Multi-organ diseases present significant challenges due to their simultaneous impact on multiple organ systems, necessitating complex and adaptive treatment strategies. Despite recent advancements in AI-powered healthcare decision support systems, existing solutions are limited to individual organ systems. They often ignore the intricate dependencies between organ system and thereby fails to provide holistic treatment recommendations that are useful in practice. We propose a novel hierarchical multi-agent reinforcement learning (HMARL) framework to address these challenges. This framework uses dedicated agents for each organ system, and model dynamic through explicit inter-agent communication channels, enabling coordinated treatment strategies across organs. Furthermore, we introduce a dual-layer state representation technique to contextualize patient conditions at various hierarchical levels, enhancing the treatment accuracy and relevance. Through extensive qualitative and quantitative evaluations in managing sepsis—a complex multi-organ disease—our approach demonstrates its ability to learn effective treatment policies that significantly improve patient survival rates. This framework marks a substantial advancement in clinical decision support systems, pioneering a comprehensive approach for multi-organ treatment recommendations.
Introduction
@@
Multi-organ diseases are characterized by the sequential or simultaneous impairment of multiple organ systems (Asim, Amin, and El-Menyar 2020). They present significant challenges in clinical management due to their complexity, the difficulty of balancing therapeutic trade-offs, and the potential for life-threatening outcomes in critically ill patients. Treating these diseases requires a holistic approach that accounts for the interdependencies between different organ systems (Tian et al. 2023). Existing guideline-based approaches treat organ systems in isolation and rely on one-size-fits-all recommendations (Whelehan, Conlon, and Ridgway 2020), which fail to address the complexities of multi-organ diseases. One example of such a disease is COVID-19, which, while primarily affecting the respiratory system, can also lead to dysfunction in the immune, nervous, and gastrointestinal systems (Thakur et al. 2021; Bhadoria and Rathore 2021). Another example is sepsis, a serious condition triggered by the body’s dysregulated response to infection. Sepsis can lead to widespread inflammation, coagulation abnormalities, and metabolic disruptions, cascading into multi-organ dysfunction, which requires comprehensive and adaptive treatment strategies (Greco et al. 2017).
Recent advances in artificial intelligence, particularly in reinforcement learning (RL), have shown promise in optimizing clinical decision-making for complex diseases. RL’s capacity to learn adaptive policies from high-dimensional, complex data makes it a powerful tool for treatment recommendations. Notably, Komorowski et al., pioneered the use of deep RL for sepsis treatment, developing a model that learns optimal treatment policies from electronic health records (EHRs) of intensive care unit (ICU) patients (Komorowski et al. 2018). Subsequent work in this area has explored new model-free approaches, such as those based on Dueling Double Deep Q-Networks (D3QN) (Raghu et al. 2017; Wu et al. 2023), as well as model-based approaches (Raghu, Komorowski, and Singh 2018). Beyond sepsis, RL has also been applied to chronic disease management, such as in the work by Zheng et al. (Zheng et al. 2021), which proposed an RL-based method for personalized diabetes and multimorbidity management by modeling health outcomes like glycemia, blood pressure, and cardiovascular disease risk. Liu et al., trained an RL model to recommend the dosage of oral antidiabetic drugs and insulin (Liu et al. 2020b). Additionally, Wang et al., proposed a model-based RL framework consisting of a patient model paired with a policy model to optimize insulin regimens through the analysis of glycemic response. (Wang et al. 2023).
@@
Despite the multi-organ nature of many major diseases, existing RL approaches have predominantly focused on recommending treatments targeting a single organ system. For example, current sepsis-targeted solutions primarily address only the cardiovascular system, neglecting other relevant organ systems (Liu et al. 2020a). This is a significant limitation, as treatments for one organ system can significantly influence the efficacy or safety of treatments for another. For instance, recommending vasopressors (VAs) to stabilize blood pressure may not be feasible in a patient with concurrent renal dysfunction, as doing so could increase renal impairment (Yagi et al. 2021). While there have been AI-based solutions developed with explicit consideration of multiple organ systems, these solutions focus on the task of diagnosis, instead of the more complex treatment recommendation task (Khan et al. 2023; Kaur and Kaur 2024).
@@
Multi-organ disease treatment recommendation tasks introduce complexities beyond that which traditional RL and non-RL-based recommendation system can effectively manage. For example, as the number of considered organ systems or treatments increase, the number of total possible combinations of actions can increase exponentially, quickly making action spaces untenable for standard single-agent RL algorithms to navigate. Additionally, patients’ measured physiological variables will relate to each organ system’s unique physiology in different ways, adding an additional layer of complexity that a multi-organ solution must account for. Consequently, there is a strong need to develop a robust and holistic treatment recommendation system tailored for multi-organ disease management.
@@
To this end, we propose a hierarchical multi-agent RL (HMARL) system. In this framework, the complex task of multi-organ treatment recommendation is divided among a hierarchy of specialized sub-agents. Each agent operates within its own localized state and action spaces, simplifying decision-making and allowing agents to focus exclusively on relevant subspaces. Carefully designed inter- and intra- agent communication mechanisms enable collaboration when necessary, alleviating the burden on individual agents and leading to more efficient training and faster convergence to optimal policies. Moreover, understanding patient states and their dynamics within localized contexts is essential for accurate treatment recommendations. For example, when treating the cardiovascular system, more focus should be on factors such as ejection fraction and cardiac enzyme levels, whereas renal treatments require attention to factors like glomerular filtration rate and electrolyte balances (Deferrari, Cipriani, and La Porta 2021). To achieve this, we propose a multi-layer hierarchical representation technique that first captures broad health indicators at the root level and then refines them into organ-specific representations, which are utilized by the corresponding agents at these levels. Collectively, this HMARL system, combined with the multi-layer hierarchical representation technique, effectively manages the complexities of multi-organ treatment recommendation.
@@
  • We introduce a compact HMARL framework that decomposes the complex task of multi-organ disease management into manageable sub-tasks, handled by specialized sub-agents operating within localized state and action spaces, both independently and collaboratively.
  • We develop a multi-layer hierarchical representation technique that learns both broad and specific patient representations, tailored to treatment context, aiding accurate decision-making at multiple levels.
@@
It must be emphasized that our solution—trained and tested on retrospective public data—serves purely as a clinical decision support system. It is not intended to take on a decision-making role, but merely offers data-driven treatment recommendations to an expert-in-the-loop in the clinical setting. There are no ethical violations in the development or use of this solution.
Methodology
Hierarchical Decomposition
Figure 1: HMARL solution architecture.
We decompose the intricate multi-organ recommendation task into a clinically meaningful hierarchy of sub-tasks, recommending seven treatments across three organ systems ( Figure 1). This structured decomposition significantly reduces the overall complexity by confining each sub-task to its own subspace, only coordinating with other subtasks when necessary.
At the top level, the root agent ($M_{Rt}$) selects from five primary actions: No-treatment ($\mathcal{A}_{Rt}^{0}$), Neuro-only ($\mathcal{A}_{Rt}^{Neu}$), Cardio-only ($\mathcal{A}_{Rt}^{Car}$), Renal only ($\mathcal{A}_{Rt}^{Ren}$) or Organ mixture ($\mathcal{A}_{Rt}^{OMix}$). Including the frequently utilized $\mathcal{A}_{Rt}^{0}$ option at the root level substantially simplifies the process by reducing the overall action space and eliminating the need to consider this non-intervention option in subsequent tasks.
Organ-specific actions ($\mathcal{A}_{Rt}^{Neu}$, $\mathcal{A}_{Rt}^{Car}$, $\mathcal{A}_{Rt}^{Ren}$) invoke respective agents ($M_{Neu}$, $M_{Car}$, $M_{Ren}$), each operating within dedicated treatment subspaces. Subsequently, $M_{Neu}$ chooses from four possible actions: S1-only ($\mathcal{A}_{Neu}^{S1}$), S2-only ($\mathcal{A}_{Neu}^{S2}$), S3-only ($\mathcal{A}_{Neu}^{S3}$), or a mixture of them ($\mathcal{A}_{Neu}^{Mix}$), each invoking a specialized sub-agent ($M_{Neu}^{S1}$, $M_{Neu}^{S2}$, $M_{Neu}^{S3}$ or $M_{Neu}^{Mix}$), recommending individual or combined treatments. $M_{Car}$ follows a similar structure, selecting from three options: IV-only ($\mathcal{A}_{Car}^{IV}$), Vaso-only ($\mathcal{A}_{Car}^{VA}$), and IV and Vaso mixture ($\mathcal{A}_{Car}^{Mix}$), invoking dedicated sub-agents ($M_{Car}^{IV}$, $M_{Car}^{VA}$ or $M_{Car}^{Mix}$), recommending individual or treatment mixtures. The recommendations for renal agent $M_{Ren}$ is comparatively simpler. Since there is no mixing between the treatments (diuretics and dialysis), $M_{Ren}$ directly chooses from four diuretic levels or. When recommending dosage mixtures for a single organ, respective sub-agents collaborate to provide a coordinated recommendation. Each treatment-specific sub-agent communicates its recommendations to the dedicated mixture agent ($M_{Neu}^{S1}$, $M_{Neu}^{S2}$, $M_{Neu}^{S3}$ to $M_{Neu}^{Mix}$ and $M_{Car}^{IV}$, $M_{Car}^{VA}$ to $M_{Car}^{Mix}$), if they were to be invoked. This additional intra-organ communication allows the mixture agent to integrate insights from respective sub-agents.
When complex multi-organ treatments are recommended ($\mathcal{A}_{Rt}^{OMix}$), the $M_{OMix}$ agent is invoked. It consults with organ-specific agents ($M_{Neu}$, $M_{Car}$, $M_{Ren}$) through inter-organ communications to access their dosage recommendations (if they were to be invoked). These auxiliary inputs help $M_{OMix}$ comprehensively understand the possible treatments and their scopes, greatly reducing its decision-making burden. Subsequently, $M_{OMix}$ invokes a combination of sub-agents (at least two pairs of $M_{OMix}^{Neu}$, $M_{OMix}^{Car}$ and $M_{OMix}^{Ren}$ agents), where these sub-agents each communicate with the opposite agents from the single-organ subtasks. For example, $M_{OMix}^{Neu}$ considers the leaf level recommendations from $M_{Car}$ and $M_{Ren}$’s sub-agents as additional information. This system benefits the agents by allowing them to adjust communicated recommendations based on their use in combination with other treatments, rather than needing to devise them from scratch. Consequently, this facilitates the agents’ ability to efficiently determine the optimal treatment combination, streamlining the decision-making process for complex, multi-organ interventions. All agents are trained exclusively on samples from their specific action subspaces. For instance, the $M_{Car}^{IV}$ agent is trained on cases where IV was the only treatment administered, while $M_{Car}^{Mix}$ agent is trained on samples involving simultaneous use of both IV and Vaso treatments.
This approach mimics a collaborative clinical setting where a lead physician oversees specialists, each prescribing treatments in their expertise to ensure comprehensive and specialized care. By distributing recommendation tasks among specific and mixture agents, our structured hierarchy efficiently handles the complexities of multi-treatment recommendations within and across organ systems. The decomposition into precise subtasks is determined by factors such as the number of treatments and their combinations. The carefully designed communication channels enables collaboration and effectively employ a divide-and-conquer strategy. We believe that the proposed hierarchical decomposition approach is both flexible and scalable, thereby making it highly suitable for various medical scenarios involving multi-organ treatments, beyond this use case.
RL Components
RL States:
@@
Accurate patient understanding is essential for effective multi-organ treatment recommendations, which require integrating complex interdependencies of physiological features and organ functions. To address this, we propose a hierarchical patient representation approach that acknowledges the varying significance of raw features at different analytical levels.
@@
At the root level, Unified State Representations are learned to extract broad health indicators and their dynamics, providing a foundational understanding of patient status. This representation is used for broader decision-making processes at the root level, and sets the stage for subsequent, more granular recommendations. At the organ levels, they are refined to learn Targeted State Representations, tailored to unique physiological requirements and and interrelationships of specific organs. For instance, in cardiac treatments, the embeddings prioritize features like ejection fraction and cardiac enzyme levels, capturing essential cardiac health indicators. In contrast, renal treatments focus on features like glomerular filtration rate and electrolyte balances, crucial for assessing renal function. This approach ensures that organ-specific recommendations are precise and relevant.
@@
This dual-layer hierarchical representation strategy balances broad applicability with detailed specificity, enhancing decision-making capabilities by considering each organ’s condition within the overall health context. The hierarchical structure comprises the following levels:
@@
Unified State Representations: Inspired by the representation learning proposed by Perera et al., (Perera, Liu, and Feng 2023) we learn the unified representations as follows. Each patient at time $t$ is represented by their raw $d$-dimensional feature, $x_{t}=\{x_{t,1},x_{t,2},\ldots,x_{t,d}\}\in\mathbb{R}$ (for details on $x_{t}$, see Appendix Section Feature Processing ). At the root agent level, these features are transformed using dense latent embeddings $E^{Rt}=\{e_{1}^{Rt},e_{2}^{Rt},\cdots,e_{d}^{Rt}\}\in\mathbb{R}^{d\times k}$. Each $k$-dimensional latent embedding vector $e_{i}^{Rt}\in\mathbb{R}^{k}$ transforms its corresponding raw feature into a more informative dense latent representation. The resultant patient-specific embeddings are represented as $F^{Rt}_{t}=\{f_{t,1}^{Rt},f_{t,2}^{Rt},\cdots,f_{t,d}^{Rt}\}\in\mathbb{R}^{d\times k}$, where $f_{t,i}^{Rt}=(x_{t,i}\cdot e_{i}^{Rt})$. These latent embeddings $e_{i}^{Rt}$ are generic at the root agent level, providing a holistic understanding of the patient by transforming each feature into a homogeneous latent space. The vector $f_{t,i}^{Rt}$ denotes the transformed representation specific to the patient at time $t$.
@@
To capture the complex interdependencies among these features, a higher-order interaction layer is introduced. This layer computes the element-wise product between all pairs of embeddings in $F_{t}^{Rt}$ resulting an interaction matrix $G_{t}^{Rt}\in\mathbb{R}^{k\times d(d+1)/2}$. Final output of the layer consists of both first- and second- order interactions, denoted by $H_{t}^{Rt}=(F_{t}^{Rt}\mid G_{t}^{Rt})$, which is then aggregated via sum pooling to generate an observation vector as $o_{t}^{Rt}=\sum_{l=1}^{d(d+3)/2}H_{t,l}^{Rt}\in\mathbb{R}^{k}$.
@@
The final core-state vector at the root level at time $t$ is constructed by concatenating the current observation vector $o_{Rt}$ with the temporal contextual state vector $c_{Rt}$, resulting in $s_{t}^{Rt}=(o_{t}^{Rt}\mid c_{t}^{Rt})\in\mathbb{R}^{2k}$. This vector is learned end-to-end during the training of the root agent, enabling the model to capture both immediate and historical physiological measurements effectively. The reduced dimensionality to $2k$ in the dense space also simplifies the complexity of subsequent decision-making processes.
@@
Targeted State Representations: The targeted (organ-level) state representations are developed by fine-tuning the generic latent embeddings $E^{Rt}$ learned at the root level to reflect the unique characteristics and critical features relevant to each organ. Accordingly, using a similar approach, we transform the raw features $x_{t}$ using specialized latent embeddings $E^{Neu}$, $E^{Car}$ and $E^{Ren}$, to obtain organ-specific state representations $s_{t}^{Neu}$, $s_{t}^{Car}$ and $s_{t}^{Ren}$. They are used as input states for training the corresponding sub-agents ($M_{Neu}$, $M_{Car}$ and $M_{Ren}$). Moreover, a concatenated targeted state representation $s_{t}^{OMix}=\left[s_{t}^{Neu}|s_{t}^{Car}|s_{t}^{Ren}\right]$ is used for training all organ mixture agents ($M_{OMix}$, $M^{Neu}_{OMix}$, $M^{Car}_{OMix}$ and $M^{Ren}_{OMix}$), ensuring they are trained on a comprehensive representation of all targeted organs.
RL Actions:

### v2

@@
Abstract.
In healthcare, multi-organ system diseases pose unique and significant challenges as they impact multiple physiological systems concurrently, demanding complex and coordinated treatment strategies. Despite recent advancements in the AI based clinical decision support systems, these solutions only focus on individual organ systems, failing to account for complex interdependencies between them. This narrow focus greatly hinders their effectiveness in recommending holistic and clinically actionable treatments in the real world setting. To address this critical gap, we propose a novel Hierarchical Multi-Agent Reinforcement Learning (HMARL) framework. Our architecture deploys specialized and dedicated agents for each organ system and facilitates inter-agent communication to enable synergistic decision-making across organ systems. Furthermore, we introduce a dual-layer state representation technique that contextualizes patient conditions at both global and organ-specific levels, improving the accuracy and relevance of treatment decisions. We evaluate our HMARL solution on the task of sepsis management, a common and critical multi-organ disease, using both qualitative and quantitative metrics. Our method learns effective, clinically aligned treatment policies that considerably improve patient survival. We believe this framework represents a significant advancement in clinical decision support systems, introducing the first RL solution explicitly designed for multi-organ treatment recommendations. Our solution moves beyond prevailing simplified, single-organ models that fall short in addressing the complexity of multi-organ diseases.
Healthcare AI, Clinical Decision Support, Dynamic Treatment Regimes, Reinforcement Learning, Personalized Medicine
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Applied computing Health informaticsccs: Information systems Decision support systems
1. Introduction
@@
Multi-organ diseases are characterized by the sequential or simultaneous impairment of multiple organs (Asim et al., 2020). They present significant challenges in clinical management due to the difficulty in balancing therapeutic trade-offs, and their potential for life-threatening outcomes. Treating these diseases requires a holistic approach that accounts for interdependencies between different organ systems (Tian et al., 2023). Existing guideline-based approaches treat organs in isolation and rely on one-size-fits-all recommendations (Whelehan et al., 2020). A recent example is COVID-19, which primarily affects the respiratory system, but could also lead to dysfunction in the immune, nervous, and gastrointestinal systems (Thakur et al., 2021; Bhadoria and Rathore, 2021). Another example is sepsis, a serious condition resulting from the body’s dysregulated response to infection. Sepsis can lead to widespread inflammation, coagulation abnormalities, and metabolic disruptions, cascading into multi-organ dysfunction (Greco et al., 2017).
@@
Recent advances in reinforcement learning (RL), have shown promise in optimizing clinical decision-making for complex diseases. Its capacity to learn adaptive policies from high-dimensional, complex data makes it a powerful tool for such tasks. Notably, deep RL has been applied to sepsis treatment, to learn dynamic treatment regimes (DTRs) from electronic health records (EHRs) of intensive care unit (ICU) patients (Komorowski et al., 2018). Subsequent work has explored model-free approaches, such as Dueling Double Deep Q-Networks (D3QN) (Raghu et al., 2017; Wu et al., 2023) and model-based approaches (Raghu et al., 2018). RL has also been applied to chronic diabetes management (Wang et al., 2023; Zheng et al., 2021; Liu et al., 2020a). Despite the multi-organ nature of many major diseases, existing RL approaches have predominantly focused on recommending treatments targeting a single organ system at a time. For example, current solutions for sepsis primarily address the cardiovascular system via fluids and vasopressor dosing optimization (Liu et al., 2020b; Komorowski et al., 2018; Raghu et al., 2017). This is a significant limitation, as treatments for one organ can significantly influence the efficacy or safety of treatments for another. For example, recommending vasopressors (VAs) to stabilize blood pressure can increase renal impairment for a patient with concurrent renal dysfunction (Yagi et al., 2021). While solutions with explicit consideration of multiple organ systems exist, their primary focus has been diagnosis, instead of complex treatment recommendation (Khan et al., 2023; Kaur and Kaur, 2024).
@@
Multi-organ disease treatment recommendation introduces complexities beyond what traditional RL and non-RL-based recommendation solutions can effectively manage. For example, as the number of organ systems or treatments increase, the number of combinations of actions increases exponentially, making action spaces untenable for standard single-agent RL algorithms to navigate. Additionally, patients’ measured physiological variables will relate to each organ system’s unique physiology in different ways, adding an additional layer of complexity that a multi-organ solution must account for. Consequently, there is a clear need to develop a robust and holistic treatment recommendation solution tailored for multi-organ system disease management.
@@
We propose a hierarchical multi-agent RL (HMARL) solution, which divides the complex task of multi-organ treatment recommendation among a hierarchy of specialized sub-agents. Each agent operates within its own localized state and action spaces, simplifying decision-making and allowing agents to focus exclusively on relevant subspaces. Carefully designed inter- and intra- agent communication mechanisms enable collaboration when necessary, alleviating the burden on individual agents and leading to more efficient training and faster convergence. Moreover, understanding patient states and their dynamics within localized contexts is essential for accurate treatment recommendations. For example, when treating the cardiovascular system, more focus should be on factors such as ejection fraction and cardiac enzyme levels, whereas renal treatments require attention to factors such as glomerular filtration rate and electrolyte balances (Deferrari et al., 2021). To achieve this, we propose a multi-layer hierarchical representation technique that first captures broad health indicators at the root level and then refines them into organ-specific representations, which are utilized by the corresponding agents at these levels. Collectively, this HMARL system, combined with the multi-layer hierarchical representation technique, effectively manages the complexities of multi-organ treatment recommendation.
@@
  • We introduce a compact HMARL framework that decomposes the complex task of multi-organ disease management into manageable subtasks, handled by specialized sub-agents operating within localized state and action spaces, independently and collaboratively.
  • We develop a multi-layer hierarchical representation technique that learns broad and specific patient representations, tailored to treatment context, aiding accurate decision-making at multiple levels.
@@
Our offline RL-based solution is trained and tested on retrospective public data as is standard in healthcare RL to ensure safety and stability. The solution serves purely as a clinical decision support system that offers data-driven treatment recommendations as an expert-in-the-loop solution. There are no ethical violations in its development or use.
2. Methodology
2.1. Hierarchical Decomposition
(a) Hierarchical decomposition and complexity reduction. The full decision space of size $\mathcal{O}(k^{N})$ is decomposed into $d$ layers, each selecting from at most $k$ options, with $b$ parallel agents per layer. Resulting complexity is reduced to $\mathcal{O}(d\cdot b\cdot k)$.
(b) Instantiation for multi-organ treatment recommendation. Root agent $M_{Rt}$ delegates to organ-specific agents or a cross-organ mixture agent $M_{OMix}$. Treatment-level decisions are made by specialized sub-agents, with intra- and inter-organ communication.
Figure 1. Hierarchical multi-agent framework and its medical instantiation. (a) A generic HMARL framework achieves scalable decision-making via depth-wise decomposition. (b) Specific application to multi-organ treatment recommendation with coordinated agent behavior.
Complex decision-making tasks such as treatment recommendation or distributed control involves complex interaction among multiple subsystems or decision domains. They often contain large combinatorial action spaces, contextual dependencies and the need for local specialization with global coordination. To address these challenges, we propose a general hierarchical MARL framework that decomposes the overall problem into multiple payers of specialized decision makers (agents). Each layer addresses a different abstraction enabling a scalable, interpretable and modular policy learning.
2.1.1. HMARL Framework Overview:
Proposed HMARL framework is organized as follows. At the highest level, a root agent selects broad high-level options or task branches. These options invoke intermediate coordinator agents that manage related subtasks within specialized domains or scopes. Leaf level sub agents are responsible for fine-grained decisions such as selecting specific actions to perform within their specialized action subspaces. Importantly, we include both intra-branch (local) and inter-branch (global) communications modules to share key information to coordinate decisions across boundaries, where necessary. This hierarchical decomposition help reduce the burden on individual agents through smaller manageable scopes and supports structured collaboration through communication pathways and shared latent representations. The framework is flexible, supporting both independent and cooperative training regimes, and scalable to complex and heterogeneous action spaces (i.e., discrete, continuous and hybrid).
2.1.2. Scalability and Complexity:
Consider a decision problem with $N$ treatment dimensions (e.g., organ or treatment modality), each with a local action space $\mathcal{A}_{i}$ of size $|\mathcal{A}_{i}|=k$. A standard flat RL agent operates on the full joint space $\mathcal{A}=\prod_{i=1}^{N}\mathcal{A}_{i}$, resulting in an exponential decision complexity of $\mathcal{O}(k^{N})$}. In contrast, proposed hierarchical MARL framework decomposes this problem across a hierarchy of depth $d$, where decisions are made sequentially by specialized agents. At each layer, an agent selects from at most $k$ options. Some agents (e.g., mixture agents) may invoke up to $b$ sub-agents in parallel to gather auxiliary recommendations. Under this setting, the per-episode decision complexity is reduced to $\mathcal{O}(d\cdot b\cdot k)$ (see Figure 1(a)). This formulation assumes conditional branching, where only relevant sub-agents are activated. If each agent (or layer) addresses on average $\overline{m}$ decision dimensions, the minimum hierarchy depth satisfies $d\geq\left\lceil\frac{N}{\overline{m}}\right\rceil,\quad\text{with}\quad b\leq\min(d,N)$.
Theoretically, $d$ could linearly scale with $N$ yet it is substantially smaller in practice. For example, in our instantiation below with $N=6$ treatment dimensions, a depth of $d=3$ is sufficient, as the decision space is structured around natural clinical groupings. Even in the worst-case scenario ($d=N$), complexity is $\mathcal{O}(N\cdot b\cdot k)$, which is polynomial and far more tractable than the exponential complexity. Moreover, each agent operates within a localized subspace $\mathcal{A}_{i}\subset\mathcal{A}$ and exclusively trained on relevant samples. This reduces gradient interference from unrelated subtasks, improves sample efficiency, and mirrors real-world collaborative clinical practice. The architecture supports plug-and-play extensibility. New treatment modalities or decision dimensions can be integrated by adding new agents without retraining the full hierarchy, allowing the overall system to scale gracefully with added complexity.
2.1.3. Clinical Use Case:
Proposed HMARL framework is domain-agnostic and can generate multiple decompositions for a given task, with varied effectivenesses based on the selected hierarchy and communication pathways. Guided by our clinical collaborators, we instantiate the framework analogous to a coordinated team of medical specialists and supervisors at practice. The model considers seven treatments across three organs (see Figure 1(b)). This decomposition restricts each subtask to a defined subspace, significantly reducing the decision making complexity while enabling selective coordination when necessary.
At the top level, the root agent $M_{Rt}$ selects one of five high-level options: No-treatment ($\mathcal{A}_{Rt}^{0}$), Neuro-only ($\mathcal{A}_{Rt}^{Neu}$), Cardio-only ($\mathcal{A}_{Rt}^{Car}$), Renal-only ($\mathcal{A}_{Rt}^{Ren}$), or Organ mixture ($\mathcal{A}_{Rt}^{OMix}$). Using the commonly selected no-treatment option at this level reduces the burden on downstream agents by pruning unnecessary branches early and reducing overall action space. When an organ-specific action is selected, the corresponding domain-level agent (i.e., $M_{Neu}$, $M_{Car}$, or $M_{Ren}$) is invoked. Neuro and Cardio agents operate similarly where each chooses between individual treatments (i.e., S1 or S2 for Neuro; IV or Vaso for Cardio) or a mixture, invoking subsequent sub-agents to recommend treatment dosages. Intra-organ communication is used for treatment mixtures ($\mathcal{A}_{Neu}^{Mix}$, $\mathcal{A}_{Car}^{Mix}$) to effectively aggregate recommendations from respective single-treatment sub-agents, enhancing local coordination. In contrast, the Renal agent $M_{Ren}$ directly selects among four diuretic levels or dialysis, as treatment mixing is not applicable. Multi-organ decisions ($\mathcal{A}_{Rt}^{OMix}$) are taken by the coordinator agent $M_{OMix}$ using the inter-organ communication to obtain their treatment recommendations, reducing the need to learn these from scratch. Accordingly, it invokes a combination of sub-agents ($M_{OMix}^{Neu}$, $M_{OMix}^{Car}$, $M_{OMix}^{Ren}$), each of which incorporates relevant signals from the other domains. For instance, $M_{OMix}^{Neu}$ considers recommendations from Cardio and Renal sub-agents. This cross-communication allows agents to adjust outputs to account for interactions, streamlining coherent multi-organ plans. Agents are trained only on samples drawn from their respective action subspaces to improve efficiency and minimize interference. For example, $M_{Car}^{IV}$ learns from IV-only cases, while $M_{Car}^{Mix}$ is trained on samples where both IV and Vaso were used.
This instantiation reflects a real-world clinical setting where a lead physician delegates to specialists. The resulting hierarchy decomposes decision-making across dedicated and mixture agents, enabling scalable inter- and intra- organ coordination. This task-specific decomposition and structured communication protocols supports modular policy learning making it highly applicable to broader multi-organ treatment scenarios.
2.2. RL Components
2.2.1. States:
@@
Accurate patient understanding is essential for effective multi-organ treatment recommendations, as it requires integrating complex interdependencies of physiological features and organ functions. Therefore, we propose a hierarchical patient representation approach that acknowledges the varying significance of raw features at different analytical levels.
@@
At the root level, Unified State Representations are learned to extract broad health indicators and their dynamics, providing a foundational understanding of patients. This representation is used for broader decision-making at the root level, and sets the stage for subsequent, more granular recommendations. At the organ levels, they are refined to learn Targeted State Representations, tailored to unique physiological requirements and interrelationships of specific organs. For instance, in cardiac treatments, the embeddings prioritize ejection fraction and cardiac enzyme levels, capturing cardiac health indicators. In contrast, renal treatments focus on glomerular filtration rate and electrolyte balances, assessing renal function. This approach ensures that organ-specific recommendations are precise and relevant.
@@
This dual-layer hierarchical representation strategy balances broad applicability with detailed specificity, enhancing decision-making considering each organ’s condition It e comprises of the following levels:
@@
Unified State Representations: Inspired by the representation learning proposed in (Perera et al., 2023) we learn the unified representations as follows. Each patient at time $t$ is represented by their raw $d$-dimensional feature, $x_{t}=\{x_{t,1},x_{t,2},\ldots,x_{t,d}\}\in\mathbb{R}$ (for details on $x_{t}$, see Appendix Section Feature Processing ). At the root agent level, these features are transformed using dense latent embeddings $E^{Rt}=\{e_{1}^{Rt},e_{2}^{Rt},\cdots,e_{d}^{Rt}\}\in\mathbb{R}^{d\times k}$. Each $k$-dimensional latent embedding vector $e_{i}^{Rt}\in\mathbb{R}^{k}$ transforms its corresponding raw feature into a more informative dense latent representation. The resultant patient-specific embeddings are represented as $F^{Rt}_{t}=\{f_{t,1}^{Rt},f_{t,2}^{Rt},\cdots,f_{t,d}^{Rt}\}\in\mathbb{R}^{d\times k}$, where $f_{t,i}^{Rt}=(x_{t,i}\cdot e_{i}^{Rt})$. These latent embeddings $e_{i}^{Rt}$ are generic at the root agent level, providing a holistic understanding of the patient by transforming each feature into a homogeneous latent space. The vector $f_{t,i}^{Rt}$ denotes the transformed representation specific to the patient at time $t$.
@@
To capture the complex interdependencies among these features, a higher-order interaction layer is introduced. This layer computes the element-wise product between all pairs of embeddings in $F_{t}^{Rt}$ resulting an interaction matrix $G_{t}^{Rt}\in\mathbb{R}^{k\times d(d+1)/2}$. Final output of the layer consists of first- and second- order interactions, denoted by $H_{t}^{Rt}=(F_{t}^{Rt}\mid G_{t}^{Rt})$, which is then aggregated via sum pooling to generate an observation vector as $o_{t}^{Rt}=\sum_{l=1}^{d(d+3)/2}H_{t,l}^{Rt}\in\mathbb{R}^{k}$.
@@
The final core-state vector at the root level at time $t$ is constructed by concatenating the current observation vector $o_{Rt}$ with the temporal contextual state vector $c_{Rt}$, resulting in $s_{t}^{Rt}=(o_{t}^{Rt}\mid c_{t}^{Rt})\in\mathbb{R}^{2k}$. This vector is learned end-to-end during the training of the root agent, enabling the model to capture immediate and historical physiological measurements effectively. The reduced dimensionality to $2k$ in the dense space also simplifies the complexity of subsequent decision-making processes.
@@
Targeted State Representations: The generic latent embeddings $E^{Rt}$ learned at the root level are fine-tuned to generate targeted (organ-level) state representations that reflect unique characteristics and critical features relevant to each organ. Accordingly, we transform the raw features $x_{t}$ using specialized latent embeddings $E^{Neu}$, $E^{Car}$ and $E^{Ren}$, to obtain organ-specific state representations $s_{t}^{Neu}$, $s_{t}^{Car}$ and $s_{t}^{Ren}$. They are used as input states for training the corresponding sub-agents ($M_{Neu}$, $M_{Car}$ and $M_{Ren}$). A concatenated targeted state representation $s_{t}^{OMix}=\left[s_{t}^{Neu}|s_{t}^{Car}|s_{t}^{Ren}\right]$ is used to train organ mixture agents ($M_{OMix}$, $M^{Neu}_{OMix}$, $M^{Car}_{OMix}$ and $M^{Ren}_{OMix}$), ensuring they are trained on a comprehensive representation of all organs.
2.2.2. Actions